# Comparison of various schemes to determine the Young’s modulus of disordered carbon nanomembranes compared to crystalline graphene

### v1

@@ Abstract
The determination of mechanical properties such as the Young’s modulus provides an important means to compare classical molecular dynamics simulations with materials. In this respect, ultra-thin materials hold several challenges: their volume is ambiguous, and different methods to determine a stress-strain relation deliver different result in particular for disordered systems. Using the example of carbon nanomembranes we show that stress-strain simulations following experimental setups deliver correct results if adjusted carefully.
@@
carbon nanomembrane; graphene; classical molecular dynamics; LAMMPS; Young’s modulus; stress-strain; barostat
@@ 1 Introduction
On the theoretical side such materials are often modelled by means of classical molecular dynamics schemes since the number of atoms needed for a trustworthy simulation is way too big for a quantum calculation for instance by means of density functional theory (DFT). Again this holds in particular for disordered, i.e., non-crystalline systems such as bio-molecular membranes or carbon nanomembranes (CNMs). The latter will be treated in the present paper. As a side remark, mechanical properties are valuable observables for such systems since due to the lack of a quantum mechanical treatment observables such a conductivity or band structure are not available.
@@
CNMs are rather soft compared to graphene; CNMs have got a modulus of about $10$ GPa [14], whereas graphene features $1000$ GPa [16, 17, 2]. The structure of CNMs is irregular and contains many holes through which for instance water permeates [18, 19]. Since CNMs constitute metastable excited states (graphite would constitute the ground state) a trustworthy simulation of mechanical properties is challenging compared to crystalline structures such as graphene or diamond.
In this paper, we discuss the determination of the Young’s modulus using three different procedures. For crystalline samples these methods yield very close results [2]. However, for disordered materials featuring both strong and weak bonds between atoms, where smallest deformations can push a system into new configurations [20], the three methods deliver rather different outcomes. Our conclusion is that both stress-strain based methods discussed in this paper are suited to calculate the elastic properties of crystalline and amorphous structures.
@@ 2.4 Stress-strain method
In order to imitate this experiment in an MD simulation, the clamping and the stretching of the material must be modeled. A static approach was chosen for this method, meaning no time integration is involved. The procedure can be used with periodic or non-periodic lateral boundary conditions, depending on the structure. In this case, p p f boundary conditions are used, i.e. the simulation box is only non-periodic in the z-direction. The method presented in Ref. [30] has been adapted for this procedure by making use of selection box regions of the size of the clamp (region command in LAMMPS). To strain the structure these regions have to move apart. Atoms in the clamp regions are excluded from the calculation of stresses by setting the forces to zero in their corresponding groups (fix setforce 0 0 0) as they are considered as rigid. To minimize the error, the clamp size is chosen to be minimal, otherwise a systematic error will occur [26]. The clamps are particularly important for non-periodic structures as a box deformation does not automatically induce an elongation of the structure. The initial length of the simulation box along strain direction is saved for the calculation of the strain at each time step. Finally, the clamps are moved outwards in discrete deformation steps using the change_box command with a predefined strain-rate. To ensure that the atoms in the clamped regions also follow the movement, the ”remap” keyword has to be specified in the change_box command.
@@ 3.2 Stress-strain method
An isotropic stress behavior is clearly visible for graphene, the mean value and its error can be combined to $E=(854\pm 5.2)$ GPa, which is in general agreement with other literature values [40]. For the carbon nanomembranes the analysis yields $E=(35.835\pm 0.835)$ GPa for CNM1 and $E=(35.705\pm 5.135)$ GPa for CNM2. As CNMs are amorphous carbon system one would expect an isotropic behavior as a stress response. This behavior is clearly visible for the CNM1-structure where the two fit functions have almost identical slopes. The offset between the two data series indicates a structure that is not fully relaxed in one spatial dimension. Care should therefore always be taken to ensure that the material is fully relaxed before the start of straining. For the CNM2-structure, two different slopes are evident indicating anisotropy in the stress behavior. The unit cell may be too small to ensure complete isotropy. Therefore, a sufficiently large structure that captures all the features of the mother material should always be used.
@@ 3.3 Barostated dynamics
To ensure that the method implemented with LAMMPS delivers correct results the Young’s modulus of graphene is calculated using the EDIP potential. A general practical advise (for every structure) is to have the structure equilibrated in the npt ensemble at desired pressure and temperature before the actual tensile experiment starts. This guarantees that there are no initial stresses or temperature differences across the sample. It is also important to monitor temperature and pressure in unstrained directions during the test to ensure that it remains constant within the elastic range, see Fig. 6.
@@
The thickness of graphene, which is input for the calculation of the modulus, is chosen to be $3.35$, compare Sec. 2.6. To avoid problems with periodic boundary effects in z-direction, the thickness of the unit cell is chosen to be much larger during the simulation. The result is then corrected with respect to the true thickness. It is important to note here that the correction factor is the ratio of the thickness of the simulation box relaxed in z-direction at the time of the start of strain and the actual thickness of the structure. The final result and its error are the averages of values in x- and y-direction and their corresponding errors. An engineering strain rate of $\varepsilon=0.1$ ps-1 and metal units with a time step of $dt=0.0001$ ps are utilized for this simulations. Simulations were carried out again at different temperatures with and without pressure control in z-direction (Fig. 7). Here a total strain of $0.02$, corresponding to 2000 simulated time steps, is considered as an elastic deformation and is therefore used to fit the linear function, see Fig. 6.
@@
Taking into account the errors, the applied temperature in a range of 50 K to 350 K does not seem to have any significant influence for the structures tested here. In general, this may of course be different. The obtained mean values of the Young’s modulus in this temperature regime are summarized in table 1 together with the values obtained by the other methods.
@@ 4 Summary
If more than a rough estimate of the Young’s modulus of crystalline 2d or quasi 2d structures, or even the tensile strength of an amorphous structure is intended, the scaling method with its calculation via the curvature of the potential energy should not be favoured. While the value for single-layer graphene is in agreement with calculations from reference [2], the moduli for CNM structures are an order of magnitude too big compared to experimental values of about 10 GPa [14] as already reported by Ehrens et al. [12]. Both stress-strain based algorithms perform much better and more reliable as they also account for other elastic deformation mechanism instead of only bond elongation.
The obtained results are much closer to the predicted 10 GPa for CNMs. Remaining discrepancies are probably due to the uncertainties of the structure itself.
Which method works best depends on the application. The clamped region stress-strain method is easier and saver to use, but only calculations at $T=0$ are possible. If it is assumed that the mechanical stability varies noticeably with temperature, this method should not be used.

### v2

@@ Abstract
The determination of mechanical properties such as the Young’s modulus provides an important means to compare classical molecular dynamics simulations with materials. In this respect, ultra-thin materials hold several challenges: their volume is ambiguous, and different methods to determine a stress-strain relation deliver different result in particular for disordered systems. Using the example of carbon nanomembranes we discuss three common approaches to the problem and show that stress-strain simulations following experimental setups deliver correct results if adjusted carefully. We provide step-by-step instructions how to perform trustworthy simulations.
@@
disordered two-dimensional carbon systems; carbon nanomembrane; graphene; classical molecular dynamics; LAMMPS; Young’s modulus; stress-strain; barostat
journal: Physica E
@@ 1 Introduction
On the theoretical side such materials are often modelled by means of classical molecular dynamics schemes since the number of atoms needed for a trustworthy simulation is way too big for a quantum calculation for instance by means of density functional theory (DFT). Again this holds in particular for disordered, i.e., non-crystalline systems such as bio-molecular membranes or carbon nanomembranes (CNMs) in contrast to regular structures such as boron nitride nanosheets on top of graphene domains that are crystalline and investigated by means of DFT in [1, 2]. In the present paper, disordered carbon nanomembranes will be treated, see Fig. 3 below. As a side remark, mechanical properties are valuable observables for such systems since due to the lack of a quantum mechanical treatment observables such a conductivity or band structure are not available. However, the technical problems addressed in this paper apply to both regular as well as irregular quasi two-dimensional systems.
@@
CNMs are rather soft compared to graphene; CNMs have got a modulus of about $10$ GPa [16], whereas graphene features $1000$ GPa [18, 19, 4]. The structure of CNMs is irregular and contains many holes through which for instance water permeates [20, 21, 22]. Since CNMs constitute metastable excited states (graphite would constitute the ground state) a trustworthy simulation of mechanical properties is challenging compared to crystalline structures such as graphene or diamond.
In this paper, we discuss the determination of the Young’s modulus using three different common procedures. For crystalline samples these methods yield very close results [4]. However, for disordered materials featuring both strong and weak bonds between atoms, where smallest deformations can push a system into new configurations [23], the three methods deliver rather different outcomes. Our conclusion is that both stress-strain based methods discussed in this paper are suited to calculate the elastic properties of crystalline and amorphous structures. For practitioner, we provide step-by-step instructions how to perform trustworthy simulations.
@@ 2.4 Stress-strain method
In order to imitate this experiment in an MD simulation, the clamping and the stretching of the material must be modeled, compare also [33, 34]. A static approach was chosen for this method, meaning no time integration is involved. The procedure can be used with periodic or non-periodic lateral boundary conditions, depending on the structure. In this case, p p f boundary conditions are used, i.e. the simulation box is only non-periodic in the z-direction. The method presented in Ref. [35] has been adapted for this procedure by making use of selection box regions of the size of the clamp (region command in LAMMPS). To strain the structure these regions have to move apart. Atoms in the clamp regions are excluded from the calculation of stresses by setting the forces to zero in their corresponding groups (fix setforce 0 0 0) as they are considered as rigid. To minimize the error, the clamp size is chosen to be minimal, otherwise a systematic error will occur [29]. The clamps are particularly important for non-periodic structures as a box deformation does not automatically induce an elongation of the structure. The initial length of the simulation box along strain direction is saved for the calculation of the strain at each time step. Finally, the clamps are moved outwards in discrete deformation steps using the change_box command with a predefined strain-rate. To ensure that the atoms in the clamped regions also follow the movement, the ”remap” keyword has to be specified in the change_box command.
@@ 3.2 Stress-strain method
An isotropic stress behavior is clearly visible for graphene, the mean value and its error can be combined to $E=(854\pm 5.2)$ GPa, which is in general agreement with other literature values [45]. For the carbon nanomembranes the analysis yields $E=(35.835\pm 0.835)$ GPa for CNM1 and $E=(35.705\pm 5.135)$ GPa for CNM2. As CNMs are amorphous carbon system one would expect an isotropic behavior as a stress response. This behavior is clearly visible for the CNM1-structure where the two fit functions have almost identical slopes. The offset between the two data series and the fact that the stress is not zero at zero strain indicates a structure that is not fully relaxed in spatial dimensions. In the present case, the offset is ignorable as both materials remain well within their linear regime, see Fig. 4 and Fig. 6. However, this is not generally the case, which is why care should always be taken to ensure that the material is (practically) fully relaxed before the start of the straining. For the CNM2-structure, two different slopes are evident indicating anisotropy in the stress behavior. The unit cell may be too small to ensure complete isotropy. Therefore, a sufficiently large structure that captures all the features of the mother material should always be used.
@@ 3.3 Barostated dynamics
To ensure that the method implemented with LAMMPS delivers correct results the Young’s modulus of graphene is calculated using the EDIP potential. A general practical advise (for every structure) is to have the structure equilibrated in the npt ensemble at desired pressure and temperature before the actual tensile experiment starts. This guarantees that there are no initial stresses or temperature differences across the sample. To equilibrate, we use a thermostat and an uncoupled barostat in all but the strained direction (in calculations with fixed z this dimension is not included), at temperature $T$ and pressure $P=1$ bar, with damping rates $\gamma_{T}=100*dt=0.01$ ps and $\gamma_{P}=10^{3}*dt=0.1$ ps for $15\cdot 10^{4}$ time steps corresponding to $150$ ps. It is also important to monitor temperature and pressure in unstrained directions during the test to ensure that they remain constant within the elastic range, see Fig. 6. This allows us to determine whether the chosen strain rate is sufficiently small to accurately represent the tensile test.
@@
The thickness of graphene, which is input for the calculation of the modulus, is chosen to be $3.35$, compare Sec. 2.6. To avoid problems with periodic boundary effects in z-direction, the thickness of the unit cell is chosen to be much larger during the simulation. The result is then corrected with respect to the true thickness. It is important to note here that the correction factor is the ratio of the thickness of the simulation box relaxed in z-direction after equilibration at the time of the start of strain and the actual thickness of the structure. Since the z-dimension also changes due to the barostat, we tested a dynamic correction factor and found that in the region we use here, neglecting this factor introduces only a negligible error compared to other error sources. The final result and its error are the averages of values in x- and y-direction and their corresponding errors. An engineering strain rate of $\varepsilon=0.1$ ps-1 and metal units with a time step of $dt=0.0001$ ps are utilized for this simulations. Simulations were carried out again at different temperatures with and without pressure control in z-direction (Fig. 7). Here a total strain of $0.02$, corresponding to 2000 simulated time steps, is considered as an elastic deformation and is therefore used to fit the linear function, see Fig. 6.
@@
Taking into account the errors, the applied temperature in a range of 50 K to 350 K does not seem to have any significant influence for the structures tested here. In general, this may of course be different. The obtained mean values of the Young’s modulus in this temperature regime are summarized in table 1 together with the values obtained by the other methods. An interesting side note is that, for CNMs, the Young’s modulus, contrary to expectations, seems to increase with temperature. Perhaps this is a material property of CNMs. Since there is currently no research on this topic, it would be interesting to further investigate in this effect. What is known from literature, is that CNMs begin the process of graphitization at temperatures of the order of $10^{3}$ K [8]. As temperature increases, some regions of the material could potentially transition towards more graphene-like structures, resulting in a higher tensile strength.
@@ 4 Summary
If more than a rough estimate of the Young’s modulus of crystalline 2d or quasi 2d structures, or even the tensile strength of an amorphous structure is intended, the popular scaling method with its calculation via the curvature of the potential energy should not be favoured. While the value for single-layer graphene is in agreement with calculations from reference [4], the moduli for disordered CNM structures are an order of magnitude too big compared to experimental values of about 10 GPa [16] as already reported by Ehrens et al. [14]. Both stress-strain based algorithms perform much better and more reliable as they also account for other elastic deformation mechanism instead of only bond elongation. The obtained results are much closer to the predicted 10 GPa for CNMs. Remaining discrepancies are probably due to the uncertainties of the structure itself.
Which of the investigated common methods works best depends on the application. The clamped region stress-strain method is easier and saver to use, but only calculations at $T=0$ are possible. If it is assumed that the mechanical stability varies noticeably with temperature, this method should not be used.
@@
In conclusion, with our investigation we provide step-by-step instructions how to perform trustworthy determinations of the Young’s modulus.